0 0
Add a comment Improve this question Transcribed image text
Answer #1

a)

P(Y = 8)

==BINOM.DIST(8,25,0.4,0)

= 0.119979

=0.12

b)

P(Y = 8)

= P(7.5 < Y < 8.5)

mean = np =25*0.4 = 10

sd =sqrt(npq) = sqrt(10*0.4*0.6) = 1.549193

hence

P(7.5 < Y < 8.5)

= 0.1132